18 June 2015: Thursday

LDAP:

  • LDAP Update: one issue in the 16th update.
    • They gave a student a new PIDM, and because adds happen before deletes… the code discovered the addition to an entry that already had a Banner persona and refused, but then went ahead and deleted that persona later.  I have manually added the new persona.
  • ACCOUNT Update: No issues in the past two days.

Backups:

  • CoM/IS informed me yesterday that med07 has been disabled since Monday.  Disabled the client and started the countdown to removal.
  • Due to the recovers and probably other actions taken in the great accidental file removal of yesterday – the H:\ disk of the windows ad environment did not back up.
  • Of course, the E:\ disk of the windowpane system is still not backing up.
  • Clones – the final job (Disk Clones) finished at 4:12 am this morning.
  • Reclaimed one (1) tape for the vault.
  • SR71943244:
    • EMC requested nsrck -L7 client.uvm.edu – it did not resolve the issue, though it did make one more backup visible.
    • I finally reached out to the networker mailing list.  Several useful responses back from them.  More reading to do on my own time tonight.

VPN:

  • No response from the last two messages on the VPN Audit.  I’ll ignore them and move along (will get them next year if needed).

Footprints:

  • Investigating parameters for auto-add customer settings for new project.

Tags:

Leave a Reply

You must be logged in to post a comment.